# TEMPO-VINE: A Multi-Temporal Sensor Fusion Dataset for Localization and Mapping in Vineyards

### v1

@@ Abstract
In recent years, precision agriculture has been introducing groundbreaking innovations in the field, with a strong focus on automation. However, research studies in robotics and autonomous navigation often rely on controlled simulations or isolated field trials. The absence of a realistic common benchmark represents a significant limitation for the diffusion of robust autonomous systems under real complex agricultural conditions. Vineyards pose significant challenges due to their dynamic nature, and they are increasingly drawing attention from both academic and industrial stakeholders interested in automation. In this context, we introduce the TEMPO-VINE dataset, a large-scale multi-temporal dataset specifically designed for evaluating sensor fusion, simultaneous localization and mapping (SLAM), and place recognition techniques within operational vineyard environments. TEMPO-VINE is the first multi-modal public dataset that brings together data from heterogeneous LiDARs of different price levels, AHRS, RTK-GPS, and cameras in real trellis and pergola vineyards, with multiple rows exceeding 100 m in length. In this work, we address a critical gap in the landscape of agricultural datasets by providing researchers with a comprehensive data collection and ground truth trajectories in different seasons, vegetation growth stages, terrain and weather conditions. The sequence paths with multiple runs and revisits will foster the development of sensor fusion, localization, mapping and place recognition solutions for agricultural fields. The dataset, the processing tools and the benchmarking results will be available at the dedicated webpage upon acceptance.
@@ I INTRODUCTION
GREENBOT [canadas2024multimodal] 9 (2 months) Tomato 10.4 km 2 1 stereo RGB - D -
MAgro [marzoa2024magro] 9 (5 months) Apple; Pear 3.06 km 1 2 stereo RGB RTK D D
ARD-VO [crocetti2023ard] 11 (3 months) Grapevine, olive.  28.8 km 1
@@
ROSARIO [soncini2024rosario] 6 (1 month) Soybean 7.3 km - stereo RGB-D RTK D D
HORTO-3DLM [barros2024spvsoap3d] 6 (3 months)
@@
0.5 km 1 2 RGB-D GPS D D
TEMPO-VINE (ours) >>12 (>>9 months)
@@
4.4 km 2 1 RGB-D RTK D D
Precision agriculture is undergoing a technological transformation, with increasing integration of robotic systems and automation aimed at improving productivity, reducing labor costs, and enabling sustainable farming practices [zhai2020decision]. Recent progress in autonomous navigation [panda2023agronav, cerrato2024deep] and environmental perception [ding2022recent] has demonstrated the potential of robotics in agricultural domain. Much of this research has focused on row-based crops, which represent roughly $75\%$ of planted fields in the United States [Bigelow:263079], addressing problems such as localization [winterhalter2021localization], path planning [salvetti2023waypoint], navigation [man2020research, martini2022], harvesting [hua2025harvesting], pruning [NAVONE2025pruning], and vegetation assessment [feng2020yield]. However, the development of reliable and generalizable solutions critically depends on the availability of high-quality datasets and standardized benchmarks to evaluate algorithmic performance in realistic settings.
The lack of datasets collected over entire seasons in operational agricultural environments that capture the complexities of the field represents a key limitation to research advancements. Robotic systems are often developed and validated in simulation [martini2023enhancing] or isolated experiments, which fail to represent the variability in terrain, vegetation, and seasonal conditions. Among the most attractive crop fields worldwide, vineyards pose significant challenges for localization and mapping due to their dynamic appearance, structural variation, and changing environmental factors throughout the year.
@@ II-A Heterogeneous LiDAR datasets
Light Detection and Ranging (LiDAR) sensors have proven highly suitable for place recognition tasks due to their invariance to illumination changes. Popular datasets, such as the Oxford [maddern20171], USyd [zhou2020developing] and NCLT [carlevaris2016university] datasets, collect data with this type of sensor during long-term campaigns to provide reliable data for developing and evaluating place recognition and simultaneous localization and mapping (SLAM) algorithms. These datasets capture LiDAR and vision information, along with seasonal changes, as they were recorded over the course of more than one year of consistent data collection sessions.
Recent studies have also focused on improving place recognition through deep learning, aiming at representations that remain robust under seasonal and appearance changes [vilella2025minkunext]. Among the most common approaches is LiDAR–vision fusion, which combines the geometric invariance of LiDAR with the rich texture information provided by cameras [komorowski2021minkloc++]. More recently, a growing trend is the fusion of different LiDAR sensors, since their varying fields of view (FoV) and coverage can produce higher-quality environmental maps [jung2025helios]. Datasets such as HeliPR [jung2024helipr] and GEODE [chen2024heterogeneous] exemplify this direction, as they integrate data collected with multiple LiDAR sensors featuring different characteristics. These datasets also provide loop-closure trajectories for SLAM evaluation. Nevertheless, they lack seasonal variability, as they were not acquired over long-term campaigns, and the captured runs are relatively short and limited to urban environments.
@@ II-B Agricultural datasets
Many agricultural datasets to date such as ARD-VO [crocetti2023ard], MAgro [marzoa2024magro] or Rosario [soncini2024rosario] primarily rely on RGB or multispectral imagery, which constrains their applicability to tasks beyond crop monitoring or visual detection. However, the lack of three-dimensional information limits progress on spatial perception problems such as mapping, navigation, and long-term place recognition. To address this gap, recent efforts like the Bacchus Long-Term (BLT) [polvara2024bacchus] or the HORTO-3DLM [barros2024spvsoap3d] datasets have started to incorporate 3D LiDAR recordings. Such developments open new opportunities for LiDAR-based place recognition (LPR) in agricultural environments. Nonetheless, those datasets only cover small fields with incomplete or costly sensor suites.
@@ III Experimental Settings
ID 01 Trellis 01, 02 Cloudy Stem, no leaves Low ($\sim 5$ cm)
12/02 Pergola 01, 02 Cloudy Stem, no leaves Low ($\sim 5$ cm)
@@
ID 05 Trellis 01, 02 Cloudy Branches, leaves Medium ($\sim 20$ cm)
06/06 Pergola 01, 02 Cloudy Branches, leaves Tall (>>40 cm)
@@
ID 10 Trellis 01, 02, 03 Cloudy Branches, leaves, fruit Medium ($\sim 20$ cm)
11/09 Pergola 01, 02 Cloudy Branches, leaves, fruit Tall (>>40 cm)
@@
Figure 4: File structure of the TEMPO-VITE dataset: the folders are organized to easily select the vineyard field, the run and the campaign date. For each experiment a bag file with metadata containing all the complete sensor data stream is provided, together with the RGB camera video and the ground truth trajectory file.
@@ IV-B Dataset Organization and Format
Sensor Trellis Pergola
@@
RGBD Camera
(RTAB-Map [rtabmap])
ATEt 8.32 5.97 4.58 1.72
@@
Velodyne VLP-16
(Fast-LIO [fastlio])
ATEt 23.93 26.46 0.58 1.54
@@
Livox MID-360
(Fast-LIO [fastlio])
ATEt 1.37 23.63 2.94 2.20
@@ V DATASET EVALUATION
In this section, we validate the data testing state-of-the-art algorithms for SLAM, both LiDAR and visual SLAM, and place recognition in the vineyard scenario. First, SLAM results are proposed to discuss the performance obtained using different sensors and the complexity with vegetation conditions in winter and summer. Finally, t
@@ V-A SLAM results
To assess the SLAM performance of the proposed dataset, two state-of-the-art ROS2-compatible algorithms were evaluated: RTAB-Map [rtabmap] for RGB-D data and Fast-LIO [fastlio] for LiDAR data. Experiments were conducted using data from campaigns 02 (March) and 07 (July), covering different seasonal conditions and two vineyard training systems: trellis and pergola. Each sensor was tested in each condition, resulting in 12 experiments in total. Performance was evaluated against ground truth using Absolute Pose Error (APE, in meters) and Relative Pose Error (RPE), with RPEt representing translational error (in meters) and RPEr rotational error (unitless). APE values exceeding 4 meters were considered failures, indicating SLAM divergence. Fast-LIO with Livox data achieved the best overall performance, with only one failure in the summer trellis scenario. With Velodyne, Fast-LIO performed well in the pergola but failed in the final rows of the trellis. In contrast, RTAB-Map consistently underperformed across all conditions, highlighting the challenges posed by the environment for RGB-D-based SLAM. Results are summarized in Table IV and Figure 5 shows the trajectories of the experiments from campaign 02.
TABLE V: Place recognition quantitative analysis: Recall@5 and Precision@5 for different thresholds in trellis vineyard (Run 01) at different seasons.
@@ V-B Place recognition results
To evaluate the place recognition capabilities of our dataset, we employed the handcrafted global descriptor Scan Context [kim2018scan]. Following an inter-session setup, we tested how well the method generalizes when revisiting the same vineyard locations at different stages of the vine growth cycle. For this evaluation, we only considered the vineyard sequences, excluding the pergola. Specifically, we focused on the Run 01 recordings, which provide the most extensive coverage of the vineyard terrain. To highlight the impact of seasonal changes on agricultural place recognition, we used the first run (recorded in February) as the database, while the subsequent monthly campaigns were used as queries in each iteration. Table V reports the quantitative results for all campaigns, using Recall@5 and Precision@5 under different positive match thresholds. The results show high performance when queries belong to the same season as the database, but performance decreases as the crop undergoes significant changes in subsequent months. This evidences the challenges posed by unstructured agricultural environments for localization. Figures 6 and 7 present the results in graphical form, showing Recall@N values and the precision–recall curves across the four months.
VI CONCLUSION
@@
The dataset will be further enriched with future campaigns to reach at least 12 months of temporal coverage. By providing multi-seasonal, multi-modal data under different vineyards and environmental conditions, TEMPO-VINE raises the bar of autonomous navigation benchmarking in agriculture, paving the way for the next generation of robotics and automation solutions.
ACKNOWLEDGEMENT
This work has been developed with the contribution of Politecnico di Torino Interdepartmental Centre for Service Robotics PIC4SeR222www.pic4ser.polito.it.

### v2

@@ Abstract
In recent years, precision agriculture has been introducing groundbreaking innovations in the field, with a strong focus on automation. However, research studies in robotics and autonomous navigation often rely on controlled simulations or isolated field trials. The absence of a realistic common benchmark represents a significant limitation for the diffusion of robust autonomous systems under real complex agricultural conditions. Vineyards pose significant challenges due to their dynamic nature, and they are increasingly drawing attention from both academic and industrial stakeholders interested in automation. In this context, we introduce the TEMPO-VINE dataset, a large-scale multi-temporal dataset specifically designed for evaluating sensor fusion, simultaneous localization and mapping (SLAM), and place recognition techniques within operational vineyard environments. TEMPO-VINE is the first multi-modal public dataset that brings together data from heterogeneous LiDARs of different price levels, AHRS, RTK-GPS, and cameras in real trellis and pergola vineyards, with multiple rows exceeding 100 m in length. In this work, we address a critical gap in the landscape of agricultural datasets by providing researchers with a comprehensive data collection and ground truth trajectories in different seasons, vegetation growth stages, terrain and weather conditions. The sequence paths with multiple runs and revisits will foster the development of sensor fusion, localization, mapping and place recognition solutions for agricultural fields. The dataset, the processing tools and the benchmarking results are available on the webpage [^1] https://sites.google.com/view/tempo-vine-dataset/dataset-overview.
@@ I INTRODUCTION
GREENBOT [3] 9 (2 months) Tomato 10.4 km 2 1 stereo RGB - D -
MAgro [20] 9 (5 months) Apple; Pear 3.06 km 1 2 stereo RGB RTK D D
ARD-VO [7] 11 (3 months) Grapevine, olive.  28.8 km 1
@@
ROSARIO [26] 6 (1 month) Soybean 7.3 km - stereo RGB-D RTK D D
HORTO-3DLM [1] 6 (3 months)
@@
0.5 km 1 2 RGB-D RTK D D
TEMPO-VINE (ours) 13 (10 months)
@@
38 km 2 1 RGB-D RTK D D
Precision agriculture is undergoing a technological transformation, with increasing integration of robotic systems and automation aimed at improving productivity, reducing labor costs, and enabling sustainable farming practices [31]. Recent progress in autonomous navigation [22, 5] and environmental perception [8] has demonstrated the potential of robotics in agricultural domain. Much of this research has focused on row-based crops, which represent roughly $75\%$ of planted fields in the United States [2], addressing problems such as localization [29], path planning[24], navigation[18], harvesting [10], pruning [21], and vegetation assessment [9]. However, the development of reliable and generalizable solutions critically depends on the availability of high-quality datasets and standardized benchmarks to evaluate algorithmic performance in realistic settings.
The lack of datasets collected over entire seasons in operational agricultural environments that capture the complexities of the field represents a key limitation to research advancements. Robotic systems are often developed and validated in simulation [19] or isolated experiments, which fail to represent the variability in terrain, vegetation, and seasonal conditions. Among the most attractive crop fields worldwide, vineyards pose significant challenges for localization and mapping due to their dynamic appearance, structural variation, and changing environmental factors throughout the year.
@@ II-A Heterogeneous LiDAR datasets
Light Detection and Ranging (LiDAR) sensors have proven highly suitable for place recognition tasks due to their invariance to illumination changes. Popular datasets, such as the Oxford [17], USyd [32] and NCLT [4] datasets, collect data with this type of sensor during long-term campaigns to provide reliable data for developing and evaluating place recognition and simultaneous localization and mapping (SLAM) algorithms. These datasets capture LiDAR and vision information, along with seasonal changes, as they were recorded over the course of more than one year of consistent data collection sessions.
Recent studies have also focused on improving place recognition through deep learning, aiming at representations that remain robust under seasonal and appearance changes [28]. Among the most common approaches is LiDAR–vision fusion, which combines the geometric invariance of LiDAR with the rich texture information provided by cameras [14]. More recently, a growing trend is the fusion of different LiDAR sensors, since their varying fields of view (FoV) and coverage can produce higher-quality environmental maps [11]. Datasets such as HeliPR [12] and GEODE [6] exemplify this direction, as they integrate data collected with multiple LiDAR sensors featuring different characteristics. These datasets also provide loop-closure trajectories for SLAM evaluation. Nevertheless, they lack seasonal variability, as they were not acquired over long-term campaigns, and the captured runs are relatively short and limited to urban environments.
@@ II-B Agricultural datasets
Many agricultural datasets to date such as ARD-VO [7], MAgro [20] or Rosario [26] primarily rely on RGB or multispectral imagery, which constrains their applicability to tasks beyond crop monitoring or visual detection. However, the lack of three-dimensional information limits progress on spatial perception problems such as mapping, navigation, and long-term place recognition. To address this gap, recent efforts like the Bacchus Long-Term (BLT) [23] or the HORTO-3DLM [1] datasets have started to incorporate 3D LiDAR recordings. Such developments open new opportunities for LiDAR-based place recognition (LPR) in agricultural environments. Nonetheless, those datasets only cover small fields with incomplete or costly sensor suites.
@@ III Experimental Settings
ID 01 Trellis 01, 02 Cloud Stem, no leaves Low ($\sim 5$ cm)
12/02 Pergola 01, 02 Cloud Stem, no leaves Low ($\sim 5$ cm)
@@
ID 05 Trellis 01, 02 Cloud Branches, leaves Medium ($\sim 20$ cm)
06/06 Pergola 01, 02 Cloud Branches, leaves Tall (>>40 cm)
@@
ID 10 Trellis 01, 02, 03 Cloud Branches, leaves, fruit Medium ($\sim 20$ cm)
11/09 Pergola 01, 02 Cloud Branches, leaves, fruit Tall (>>40 cm)
@@
Figure 4: File structure of the TEMPO-VINE dataset: the folders are organized to easily select the vineyard field, the run and the campaign date. For each experiment a bag file with metadata containing all the complete sensor data stream is provided, together with the RGB camera video and the ground truth trajectory file.
@@ IV-B Dataset Organization and Format
Sensor RMSE Trellis Pergola
@@
RGBD Camera (RTAB-Map [16]) ATEt 8.32 5.97 4.58 1.72
@@
Velodyne VLP-16 (Fast-LIO [30]) ATEt 23.93 26.46 0.58 1.54
@@
Velodyne VLP-16 (LIO-SAM [25]) ATEt 1.93 3.30 3.78 1.09
RPEr 0.45 0.57 0.94 0.74
RPEt 0.03 0.03 0.04 0.03
Livox MID-360 (Fast-LIO [30]) ATEt 1.37 23.63 2.94 2.20
@@ V DATASET EVALUATION
In this section, we validate the data testing state-of-the-art algorithms for SLAM, both LiDAR and visual SLAM, and place recognition in the vineyard scenario. The results demonstrate the performance obtained using different sensors and the complexity generated from vegetation conditions in winter and summer.
@@ V-A SLAM results
To assess the SLAM performance of the proposed dataset, three state-of-the-art ROS2-compatible algorithms were evaluated: RTAB-Map [16] for RGB-D data, Fast-LIO [30] and LIO-SAM [25] for LiDAR data. Experiments were conducted using data from campaigns 02 (March) and 07 (July), covering different seasonal conditions and two vineyard training systems: trellis and pergola. Each sensor was tested in each condition, resulting in 12 experiments in total. Performance was evaluated against ground truth using Absolute Pose Error (APE, in meters) and Relative Pose Error (RPE), with RPEt representing translational error (in meters) and RPEr rotational error (unitless). APE values exceeding 4 meters were considered failures, indicating SLAM divergence. LIO-SAM obtained the most robust result overall, performing successfully in all seasons. However, the algorithm is compatible only with ring-organized data like Velodyne. Fast-LIO with Livox data achieved precise performace, only failing in the summer trellis scenario. With Velodyne, Fast-LIO performed well in the pergola but failed in the final rows of the trellis. In contrast, RTAB-Map consistently underperformed across all conditions, highlighting the challenges posed by the environment for RGB-D-based SLAM. Results are summarized in Table IV and Figure 5 shows the trajectories of the experiments from campaign 02.
TABLE V: Place recognition quantitative analysis with Scan Context[13]: Recall@5 and Precision@5 for different thresholds and seasons in trellis vineyard (Run 01).
@@
TABLE VI: Place recognition quantitative analysis of learned methods: Recall@1% and Recall@1.
Sensor Camp Backbone Recall@1% Recall@1
VLP-16 02-March PointNetVLAD 0.68 -
02-March MinkLoc3Dvs 0.42 0.32
04-May PointNetVLAD 0.52 -
04-May MinkLoc3Dv2 0.49 0.30
06-June PointNetVLAD 0.47 -
06-June MinkLoc3Dv2 0.38 0.30
Mid-360 02-March PointNetVLAD 0.59 -
02-March MinkLoc3Dv2 0.39 0.32
04-May PointNetVLAD 0.46 -
04-May MinkLoc3Dv2 0.35 0.30
06-June PointNetVLAD 0.50 -
06-June MinkLoc3Dv2 0.40 0.29
@@ V-B Place recognition results
To evaluate the place recognition capabilities of our dataset, we used a handcrafted global descriptor, Scan Context [13], and two state-of-the-art deep learning global descriptors, PointNetVLAD [27] and MinkLoc3Dv2 [15]. Following an inter-session setup, we tested how well the method generalizes when revisiting the same vineyard locations at different stages of the vine growth cycle. For this evaluation, we only considered the vineyard sequences, excluding the pergola. Specifically, we focused on the Run 01 recordings, which provide the most extensive coverage of the vineyard terrain. To highlight the impact of seasonal changes on agricultural place recognition, we used the first run (recorded in February) as the database, while the subsequent monthly campaigns were used as queries in each iteration. Table V reports the Scan Context quantitative results for all campaigns, using Recall@5 and Precision@5 under different positive match thresholds. On the other hand, Table VI shows the results for the selected months with the two learned descriptors in terms of Recall@1% and Recall@1, using Run 01 from the selected campaign as the training set and Run 02 as the evaluation set. A correct match is selected in a maximum range of 5 meters.
The results show high performance when queries belong to the same season database, but performance decreases as the crop undergoes significant changes in subsequent months. Furthermore, the geometry-based descriptor Scan Context outperforms the two learned methods, which were designed for urban environments. This evidences the challenges posed by agricultural environments for localization. Figures 6 and 7 present the results in graphical form, showing Recall@N values and the precision–recall curves across the different months.
VI CONCLUSIONS
@@
As future work, the dataset will be extended to reach two years temporal coverage, including all the four seasons, and different trajectories. SLAM benchmarking will be enriched and open to the community, carrying out introspection analyses about strengths and failures of methods over the different field conditions. By providing multi-seasonal, multi-modal data under different vineyards and environmental conditions, TEMPO-VINE raises the bar of autonomous navigation benchmarking in agriculture, paving the way for the next generation of robotics and automation solutions.